OpenStudy (anonymous):

write the decimal as a fraction in simplest form -0.3

OpenStudy (anonymous):

-3/10

OpenStudy (anonymous):

\[\frac{-0.3}{1}*\frac{10}{10}=\frac{-3}{10}\]